OpenStudy (anonymous):

the speed of a passenger train is 12mph faster than the speed of a freight train. The passenger train travels 350 miles in the same time it takes the freight train to travel 270 miles. Find the speed of each train

OpenStudy (dumbcow):

d = r*t passenger: 350 = (r+12)t = rt +12t freight: 270 = rt use substitution 350 = 270 + 12t 80 = 12t 20/3 = t 270 = (20/3)r 40.5 = r

OpenStudy (anonymous):

answer was incorrect